# Forecasting Treatment Response with Deep Pharmacokinetic Encoders

### v7

@@ Abstract
Forecasting healthcare time series data is vital for early detection of adverse outcomes and patient monitoring. However, forecasting is challenging in practice due to variable medication administration and unique pharmacokinetic (PK) properties for each patient. To address these challenges, we propose a novel hybrid global-local architecture and a PK encoder that informs deep learning models of patient-specific treatment effects. We showcase the efficacy of our approach in achieving significant accuracy gains for a blood glucose forecasting task using both realistically simulated and real-world data. Our hybrid global-local architecture improves over patient-specific models by 15.8%  on average. Additionally, our PK encoder surpasses baselines by up to 16.4% on simulated data and 4.9% on real-world data for individual patients during critical events of severely high and low glucose levels. Our code is available at https://github.com/PotosnakW/neuralforecast/tree/pk_paper_code.
keywords:
Multi-Input Forecasting, Healthcare, Pharmacokinetics, Deep Learning
@@ 1 Introduction
Leveraging the predictive power of exogenous variables presents unique challenges, especially in healthcare, where the frequency of patient signals and exogenous factors, such as medication doses or treatments, often exhibit a mismatch in temporal resolution. Furthermore, while the medication dose is usually known, one cannot directly observe the plasma concentration of the drug, as illustrated in Fig. 1. Consequently, the impact of dynamic changes in treatment resulting from drug absorption and elimination remains hidden.
Figure 1: Insulin dose is usually known and recorded in data, whereas plasma insulin concentration is unobservable and varies over time.
We consider the blood glucose forecasting task while taking into account latent, time-dependent treatment effects. Patients with diabetes must regularly monitor and regulate their blood glucose values to prevent hypoglycemic and hyperglycemic events, which can cause irreversible organ damage or, in severe cases, be fatal. Hyperglycemic and hypoglycemic episodes are responsible for numerous hospital admissions each year. A 2009 report on hospital admissions for hyperglycemia indicates over 500,000 patient-days and costs of \$2.4 billion each year for hyperglycemia alone (Kitabchi et al., 2009). Another report on hypoglycemic hospital admissions suggests a stable rate of 5.6-7.3 admissions per 1,000 patients annually between 2016-2019 (Pilla et al., 2021). With over 38 million patients with Diabetes Mellitus in the United States alone, this equates to a conservative estimate of 212,800 patients hospitalized for hypoglycemic episodes each year. Capturing the dynamic effects of insulin is crucial for accurately predicting its impact on blood glucose levels to anticipate and proactively address abnormal levels to prevent adverse patient events.
@@
  • Multiple Dose Effects Framework. We provide a computationally efficient framework for encoding the effects of multiple doses. We verify this approach mathematically under the assumption of linear PK.
  • State-of-the-Art Results. We obtain significant accuracy improvements on large scale simulated and real-world patient blood glucose forecasting datasets.
@@ 1.1 Pharmacokinetics Background
PK modeling is a natural solution to the aforementioned problem, as it can unveil medication effects through knowledge of drug movement within the body. For example, PK provides a well-established understanding of plasma insulin concentration changes over time based on drug absorption and elimination characteristics  (Binder et al., 1984). Empirical and invasive studies show that concentration-time profiles for subcutaneous injections follow right-skewed curves. The degree of skew varies with insulin type and patient-specific factors, as shown in Fig. 2, where the area under the concentration curve measures drug bioavailability, or the fraction of the absorbed dose that reaches the intended site of systemic circulation intact. Most drugs used in clinical practice are assumed to follow linear PK (Rosenbaum, 2017). Under this model, the area under the concentration curve is directly proportional to the drug dose. A complete reference for supporting PK information regarding concentration-time profiles, insulin analogs, bioavailability, first-order kinetics, and linear PK can be found in Appendix B.
@@ 2 Related Work
Blood Glucose Forecasting. Neural Forecasting has previously been applied to the field of blood glucose forecasting, outperforming the accuracy of classic statistical methods. RNNs/LSTMs have been employed both as models and baselines in numerous studies (Qian et al., 2021; Li et al., 2019a; Fox et al., 2018; Rabby et al., 2021; Rubin-Falcone et al., 2020, 2022; Miller et al., 2020). Similarly, CNNs (Xie and Wang, 2020; Li et al., 2019a, b) and MLPs (McShinsky and Marshall, 2020; Miller et al., 2020; Jahangir et al., 2017) have been utilized in proposed models and as baselines.
Blood Glucose Treatment Effect Modeling. Previous studies have successfully modeled treatment effects on patient blood glucose levels. Miller et al. present one of the first hybrid statistical-and-physiological model that combines glucose-insulin dynamics and deep learning (Miller et al., 2020). They present the Deep T1D Simulator (DTD-Sim) model, which uses a deep state-space model to describe parameters representing insulin and carbohydrate absorption rates and sensitivities of the UVA/Padova simulator that simulates blood glucose levels. Odnoblyudova et al. demonstrated enhanced accuracy in predicting blood glucose levels using Multi-output Gaussian Processes (MOGPs) to model composite effects of carbohydrate and fat components from meals. (Odnoblyudova et al., 2023).
Despite recent advancements in forecasting treatment effects, significant challenges remain, including:
Sparse data.
Limited treatment data can hinder deep learning models from accurately learning treatment responses (Rubin-Falcone et al., 2022; Hussain et al., 2021). To address this challenge, we integrate domain knowledge into the deep learning models, allowing them to more effectively encode the time-dependent effects of these variables and improve forecasting performance.
Composite treatment effects.
Unlike meals that consist of various components such as carbohydrates and fats, insulin is a homogenous protein hormone composed of a single type of molecule. Therefore, modeling its component effects is not applicable (Odnoblyudova et al., 2023). Instead, we aim to model patient-specific holistic effects of the drug as well as composite effects of multiple doses.
Inflexible models.
Previous glucose forecasting models, like the UVA/Padova simulator, are specific to blood glucose dynamics and inflexible to other forecasting targets and variables (Visentin et al., 2016; Miller et al., 2020). Our approach leverages a deep learning model to generate predictions. This distinction provides greater flexibility, as our forecasting model is not constrained by a specific physiological model and can accommodate various deep learning models and forecasting targets and inputs.
Model application scope.
Prior work that integrates PK domain knowledge in deep learning models has been used in the context of predicting disease progression for cancer and COVID-19 with long-term horizons of interest (Hussain et al., 2021; Qian et al., 2021). In contrast, we leverage linear PK assumptions to create a framework that can be directly applied to various medications used in clinical practice and is easily adapted to other treatments.
Individualized networks.
Studies often use individualized networks, limiting cross-learning. Our novel global-local architecture allows efficient training of a more accurate global model applicable to multiple patients.
@@ 3 Methods
We aim to forecast blood glucose levels 30 minutes into the future consistent with prior work (Rubin-Falcone et al., 2022; Xie and Wang, 2020; Li et al., 2019b), based on a 10 hour history of blood glucose, carbohydrate (CHO), insulin basal, and insulin bolus values. We choose a 10-hour duration to balance sufficient historical information with GPU memory constraints. the duration of the rapid-acting insulin (Humalog and Novolog) used in insulin pumps by patients in our cohort typically ranges from 2 to 5 hours (ins, 2018, ). Therefore, the selection of a 10-hour lag time can fully cover multiple dose durations, allowing the model to learn treatment effects across various dosages and durations.
@@ 3.1 PK Encoder
Medication administrations are represented as instantaneous, sparse variables in time series data. We propose a novel PK encoder that informs deep learning models of time-dependent plasma drug concentration to enable more accurate forecasting of treatment outcomes as shown in Fig. 3. The encoder consists of a function, $C$, that takes in sparse exogenous variables and generates a concentration-time profile for a specified dose, $\textbf{x}^{(i)}$, using the following equation:
@@
where $t$ is the time and $\textbf{k}^{(i)}$ is a PK parameter affecting elimination rate that specific to patient, $i$. Here, $C$ is equivalent to the log-normal probability density function (pdf) with $\sigma=\textbf{k}$ and $\mu=1$ and a scaling factor, x. The intuition for the log-normal shape of the concentration curves stems from pharmacokinetics literature as outlined in Section 1.1 and Appendix B.
@@
Under linear PK assumptions, PK parameters, including bioavailability, are assumed to be constant with respect to the drug. Given this assumption, we show mathematically that the area under the concentration curve is bounded by the dose in Proposition A.1 in Appendix A.5. To model bioavailability and incorporate dose information directly into the concentration curves, we use the dose, x, as a scaling factor in $C$. This ensures that the area under the curve integrates to the dose value, under the simplifying assumption that all insulin administrations adhere to 100% bioavailability, which we show is upper bounded by the dose. Scaling the concentration curve changes its magnitude but does not alter the PK parameters or shape of the curve.
@@
The encoder is trained end-to-end with the model. During model training, sparse insulin features are passed into the PK encoder along with their respective $\textbf{k}^{(i)}$ parameters, as shown in Fig. 4. The encoder than replaces sparse medication features in the training window with the corresponding concentration curve. For a single dose, x, of insulin administered at time, $t$, we obtain the concentration, $\widecheck{\textbf{x}}^{(i)}$, at that time for patient, $i$:
@@
For each time step within the forecast horizon, $H$, forecasts, $\hat{\textbf{y}}$, are generated by the model, $f_{\theta}$, as a function of static features, s, treatment concentration curves, $\widecheck{\textbf{x}}$, and historical glucose values y recorded in the time prior to $t$ based on the lag time, $L$,
$\hat{\textbf{y}}^{(i)}_{t:t+H}$ $=f_{\theta}(\textbf{y}^{(i)}_{t-L:t},\widecheck{\textbf{x}}^{(i)}_{t-L:t},\textbf{s}^{(i)}).$ (3)
In the proposed context, s includes patient age, weight, one-hot encoded subject identification number, and insulin pump type. $\widecheck{\textbf{x}}_{t-L:t}$ includes concentration curves for basal insulin, $\widecheck{\textbf{x}}_{\text{basal}}$, bolus insulin, $\widecheck{\textbf{x}}_{\text{bolus}}$, and CHO, $\widecheck{\textbf{x}}_{\text{CHO}}$. We include plasma treatment effects of CHO because, similar to drugs, it undergoes absorption and metabolism processes, resulting in glucose and other sugars entering the bloodstream.
The PK encoder training framework is outlined in Appendix A.6, which can be integrated into any deep learning architecture capable of handling exogenous variables alongside modeling the target variable.
@@ 3.2 Modeling Multiple Dose Effects
Multiple insulin doses may occur at close intervals in an event often referred to as “insulin stacking” as shown in Fig. 11. To model PK effects for multiple doses in an input window, we generate a concentration curve for each dose event and sum concentration curves across the windows. Under the assumption of linear pharmacokinetics (outlined in 1.1 and Appendix B), we show that the overall effect, or concentration, of $n$ doses of the same drug is equal to the sum of the individual effects which is verified mathematically in Proposition A.2 in Appendix A.5. To perform vectorized operations that enable generating concentration curves for multiple doses in an input window, the PK encoder leverages a matrix, $\textbf{w}\in\mathbb{R}^{L\times L}$, where $L$ refers to the lag time. This approach allows for efficient modeling of multiple effects in $\mathcal{O}(1)$ time. Each row of w represents a time count up to the input time where the count shifts one time step for each subsequent row. Concentration curves are generated for each row of w using the dose at the corresponding time index. If no dose is administered at a particular time, $t$, the corresponding row of w indexed by $t$ will be zeroed out. The final concentration curve feature, which can reflect the impact of multiple doses, is generated by aggregating the individual concentration curves across time steps:
@@ 3.3 Modeling Dose-Dependent Effects
CGMs administer the same type of insulin (rapid-acting) for basal and bolus doses. However, bolus insulin doses are generally much larger than basal doses as highlighted in Figs. 9 and 10 in Appendix A.1. To account for potential dose-dependent effects, we treat each dose range ($\textbf{k}_{\text{bolus}}$ and $\textbf{k}_{\text{basal}}$) as a separate entity. By treating each dose range as a separate entity, we can capture potential PK variations without resorting to a fully nonlinear PK model. This method simplifies the modeling process by maintaining linear assumptions within each dose category. It avoids the complexity of a full nonlinear model while still accounting for potential differences between the two categories of low-dose (basal) and high-dose (bolus) insulin kinetics. Specifically, we estimate $\textbf{k}_{\text{basal}}$ and $\textbf{k}_{\text{bolus}}$ values. For a single dose of basal insulin, $\textbf{x}_{\text{basal}}$, or bolus insulin, $\textbf{x}_{\text{bolus}}$, administered at time, $t$, we obtain the plasma insulin concentration at that time for patient, $i$:
@@
Treating each dose as a different entity is an approximation and may not capture all nuances of dose-dependent kinetics. All physiological models are approximations of complex biological processes. Modeling dose-dependent effects in a more granular manner could potentially provide more precise results if such effects are present. However, our aim is to develop minimal models validated under established linear PK assumptions from the pharmacology literature, striking a balance between simplicity and accuracy. Our framework ensures broader applicability across various medications where a majority of drugs are assumed to adhere to linear PK (Rosenbaum, 2017).
The equations presented in the paper are shown in a patient-specific manner to highlight that the PK encoder generates individualized information. However, these equations can be applied to produce patient-specific outputs for multiple patients simultaneously using vectorized operations.

### v8

# Global Deep Forecasting with Patient-Specific Pharmacokinetics

@@ Abstract
Forecasting healthcare time series data is vital for early detection of adverse outcomes and patient monitoring. However, it can be challenging in practice due to variable medication administration and unique pharmacokinetic (PK) properties of each patient. To address these challenges, we propose a novel hybrid global-local architecture and a PK encoder that informs deep learning models of patient-specific treatment effects. We showcase the efficacy of our approach in achieving significant accuracy gains in a blood glucose forecasting task using both realistically simulated and real-world data. Our PK encoder surpasses baselines by up to 16.4% on simulated data and 4.9% on real-world data for individual patients during critical events of severely high and low glucose levels. Furthermore, our proposed hybrid global-local architecture outperforms patient-specific PK models by 15.8%, on average. Our code is available at https://github.com/PotosnakW/neuralforecast/tree/pk_paper_code
@@ 1 Introduction
Leveraging the predictive power of exogenous variables presents unique challenges, especially in healthcare, where the frequency of patient signals and exogenous factors, such as medication doses or treatments, often exhibit a mismatch in temporal resolution. Furthermore, while the drug dose can be observed and recorded, the body’s interaction with the drug cannot be directly observed and is typically inferred from its plasma concentration, as illustrated in Fig. 1. Consequently, the recorded drug dose alone cannot capture time-dependent drug kinetics, including processes such as absorption and elimination.
Figure 1: Insulin dose is recorded, but plasma insulin concentration is not directly measurable and typically requires invasive procedures and laboratory tests.
Consider the task of blood glucose monitoring for patients with diabetes who must regularly monitor and regulate their blood glucose level. Careful and prompt administration of insulin and intake of carbohydrates is required to prevent hyper- and hypoglycemic events, which can cause irreversible organ damage or, in severe cases, be fatal. Hyperglycemic and hypoglycemic episodes are responsible for numerous hospital admissions each year. A 2009 report on hospital admissions indicates over 500,000 patient-days and costs of \$2.4 billion each year for hyperglycemia alone (Kitabchi et al., 2009). Another report on hypoglycemic hospitalizations suggests a stable rate of 5.6-7.3 admissions per 1,000 patients annually between 2016-2019 (Pilla et al., 2021). With over 38 million patients with Diabetes Mellitus in the United States alone, this equates to a conservative estimate of 212,800 patients hospitalized for hypoglycemic episodes each year.
For healthcare challenges like this, advanced forecasting technologies can help patients and clinicians monitor, anticipate, and proactively address emerging abnormalities to prevent adverse events. However, accurate forecasting requires models that can leverage exogenous information, such as treatments, while capturing latent treatment kinetics and their effects to generate reliable forecast trajectories. To address this, we propose a hybrid model that integrates cohort-level information with patient-specific pharmacokinetic characteristics to improve the predictability of blood glucose responses to treatment.
@@
  3. (iii)
    Multiple Dose Effects Framework. We provide a computationally efficient framework for encoding the effects of multiple doses. We verify this approach mathematically under the assumption of linear pharmacokinetics.
  4. (iv)
    State-of-the-Art Results. We obtain significant accuracy improvements on large-scale simulated and real-world patient blood glucose forecasting datasets.
@@ 1.1 Pharmacokinetics Background
PK modeling is a natural solution to the aforementioned problem, as it can reveal the effects of the medication over time through knowledge of drug movement within the body. For example, PK provides a well-established understanding of the changes in plasma insulin concentration over time based on drug absorption and elimination characteristics  (Binder et al., 1984). Empirical and invasive studies show that concentration-time profiles for subcutaneous injections follow right-skewed curves. The extent of skew varies with insulin type and patient-specific factors, as shown in Fig. 2, where the area under the concentration curve measures drug bioavailability, or the fraction of the absorbed dose that reaches the intended site of systemic circulation intact. Most drugs used in clinical practice are assumed to follow linear PK (Rosenbaum, 2017). Under this model, the area under the concentration curve is directly proportional to the drug dose. A complete reference for supporting PK information regarding concentration-time profiles, insulin analogs, bioavailability, first-order kinetics, and linear PK can be found in Appendix B.
@@ 2 Related Work
Blood Glucose Forecasting. Neural Forecasting has previously been used to predict blood glucose levels, outperforming traditional statistical methods. RNNs and LSTMs have been used both as models and as baselines in numerous studies (Qian et al., 2021; Li et al., 2019a; Fox et al., 2018; Rabby et al., 2021; Rubin-Falcone et al., 2020, 2022; Miller et al., 2020). Similarly, CNNs (Xie and Wang, 2020; Li et al., 2019a, b) and MLPs (McShinsky and Marshall, 2020; Miller et al., 2020; Jahangir et al., 2017) have been used in proposed models and as baselines.
Modeling Blood Glucose Treatment Effects.
Prior works have looked at modeling treatment effects on blood glucose levels. Miller et al. introduced one of the first hybrid statistical-physiological models, the Deep T1D Simulator (DTD-Sim), which integrates glucose-insulin dynamics with deep learning through a deep state-space model (Miller et al., 2020). Odnoblyudova et al. improved the prediction of blood glucose using multi-output Gaussian processes (MOGP) to model the composite effects of meal components, such as carbohydrates and fats (Odnoblyudova et al., 2023). Unlike meals, insulin is a homogeneous protein hormone, making component-based modeling inapplicable to our task (Odnoblyudova et al., 2023). Instead, our objective is to capture the holistic patient-specific effects of the drug and the composite effects of multiple doses.
Despite advances in treatment effect modeling, current approaches face limitations, including (1) reliance on individualized, subject-specific models and (2) inflexible architectures that hinder adaptability to various base models, forecasting tasks, and model inputs. The next paragraph outlines these limitations and how our proposed contributions address them.
Previous work has developed individualized networks that overlook the benefits of developing global models that learn from various, diverse data series. Such cross-learning on diverse datasets can enhance generalization abilities, as demonstrated with Time Series Foundation Models (TSFMs) (Goswami et al., 2024). Our hybrid global-local architecture enables cohort-level training to develop a global model applicable to multiple patients while also preserving patient-specific information for personalized forecasts. Regarding architectures, models such as the UVA / Padova simulator depend on predefined ODEs for glucose dynamics, restricting adaptability to other targets and exogenous inputs (Visentin et al., 2016; Miller et al., 2020). Similarly, deep learning approaches for sparse exogenous variables often rely on fixed base models like LSTMs (Rubin-Falcone et al., 2022). Our proposed contributions support diverse deep learning models, and are flexible to forecasting targets and input variables, ensuring future adaptability as more advanced tools emerge. Other prior work that proposes the integration of PK knowledge with deep learning focuses on domains such as long-term disease progression in cancer and COVID-19, where such PK models are not directly applicable to blood glucose and insulin kinetics. (Hussain et al., 2021; Qian et al., 2021). Our study provides a formal framework for integrating PK in deep learning under linear PK assumptions that are applicable to many drugs used in clinical practice (Rosenbaum, 2017).
@@ 3 Methods
We aim to forecast blood glucose levels 30 minutes into the future consistent with prior work (Rubin-Falcone et al., 2022; Xie and Wang, 2020; Li et al., 2019b), based on a 10 hour history of blood glucose, carbohydrate (CHO), insulin basal, and insulin bolus values. We choose a 10-hour duration to balance sufficient historical information with our GPU memory constraints. Moreover, the duration of rapid-acting insulin (Humalog and Novolog) used in insulin pumps by patients in our cohort typically ranges from 2 to 5 hours (Clinic, 2018; Association, n.d.). Thus, a 10-hour timespan fully captures multiple dose durations, enabling the model to learn treatment effects across various dosages and durations.
@@ 3.1 PK Encoder
Medication administrations are represented as instantaneous sparse variables in time series data. We propose a novel PK encoder that informs deep learning models of time-dependent plasma drug concentration to enable more accurate forecasting of treatment outcomes as shown in Fig. 3. The encoder consists of a function, $C$, that takes in sparse exogenous variables and generates a concentration-time profile for a specified dose, $\textbf{x}^{(i)}$, using the following equation:
@@
where $t$ is the time and $\textbf{k}^{(i)}$ is a PK parameter affecting elimination rate specific to the patient, $i$. Here, $C$ is equivalent to the log-normal probability density function (pdf) with $\sigma=\textbf{k}$ and $\mu=1$ and a scaling factor, x. The intuition for the log-normal shape of the concentration curves stems from PK literature, as outlined in Section 1.1 and Appendix B.
@@
Under linear PK assumptions, the PK parameters, including bioavailability, are assumed to be constant with respect to the drug. Given this assumption, we show mathematically that the area under the concentration curve is bounded by the dose in Proposition A.1 in Appendix A.4. To model bioavailability and incorporate dose information directly into the concentration curves, we use the dose x, as a scaling factor in $C$. This ensures that the area under the curve integrates to the dose value, under the simplifying assumption that all insulin administrations adhere to 100% bioavailability, which we show is upper bounded by the dose. Scaling the concentration curve changes its magnitude, but does not alter the PK parameters or shape of the curve.
@@
The encoder is trained end-to-end with the model. During training, sparse insulin features are passed into the PK encoder along with their respective $\textbf{k}^{(i)}$ parameters, as shown in Fig. 4. The encoder then replaces sparse medication features in the training window with the corresponding concentration curves. For a single dose, x, of insulin administered at time, $t$, we obtain the concentration, $\widecheck{\textbf{x}}^{(i)}$, at that time for patient, $i$:
@@
For each time step within the forecast horizon, $H$, forecasts, $\hat{\textbf{y}}$, are generated by the model, $f_{\theta}$, as a function of static features, s, treatment concentration curves, $\widecheck{\textbf{x}}$, and historical glucose values y recorded up to and including time $t$ based on the lag time, $L$,
$\hat{\textbf{y}}^{(i)}_{t+1:t+H}$ $=f_{\theta}(\textbf{y}^{(i)}_{t-L:t},\widecheck{\textbf{x}}^{(i)}_{t-L:t},\textbf{s}^{(i)}).$ (3)
In the proposed context, s includes patient age, weight, one-hot encoded subject identification number, and insulin pump type. $\widecheck{\textbf{x}}_{t-L:t}$ includes concentration curves for basal insulin, $\widecheck{\textbf{x}}_{\text{basal}}$, bolus insulin, $\widecheck{\textbf{x}}_{\text{bolus}}$, and CHO, $\widecheck{\textbf{x}}_{\text{CHO}}$. We include treatment effects of carbohydrates because, similar to drugs, they undergo absorption and metabolism processes, resulting in glucose and other sugars entering the bloodstream.
@@ 3.2 Modeling Multiple Dose Effects
Multiple insulin doses may occur at close intervals in an event often referred to as “insulin stacking” as shown in Fig. 3 and Fig. 14. To model PK effects for multiple doses in an input window, we generate a concentration curve for each dose event and sum concentration curves across the windows. Under the assumption of linear pharmacokinetics (outlined in 1.1 and Appendix B), we show that the overall effect, or concentration, of $n$ doses of the same drug is equal to the sum of the individual effects which is verified mathematically in Proposition A.2 in Appendix A.4. To perform vectorized operations that enable generating concentration curves for multiple doses in an input window, the PK encoder leverages a matrix, $\textbf{w}\in\mathbb{R}^{L\times L}$, where $L$ refers to the lag time. This approach allows for efficient modeling of multiple effects in $\mathcal{O}(1)$ time. Each row of w represents a time count up to the input time where the count shifts one time step for each subsequent row. Concentration curves are generated for each row of w using the dose at the corresponding time index. If no dose is administered at a particular time, $t$, the corresponding row of w indexed by $t$ will be zeroed out. The final concentration curve feature, which can reflect the impact of multiple doses, is generated by aggregating the individual concentration curves across time steps:
@@
The PK encoder training algorithm is described in Appendix A.5, which can be integrated into any deep learning architecture capable of handling exogenous variables in addition to modeling the target variable.
@@ 3.3 Modeling Dose-Dependent Effects
Insulin pumps administer the same type of insulin (rapid-acting) for the basal and bolus doses. However, bolus insulin doses are generally much larger than basal doses as highlighted in Figs. 12 and 13 in Appendix A.1. To account for potential dose-dependent effects, we treat each dose range ($\textbf{k}_{\text{bolus}}$ and $\textbf{k}_{\text{basal}}$) as a separate entity, so that we can capture potential PK variations without resorting to a fully nonlinear PK model. This method simplifies the modeling process by maintaining linear assumptions within each dose category. It avoids the complexity while still accounting for potential differences between the two categories of low-dose (basal) and high-dose (bolus) insulin kinetics. Specifically, we separately estimate $\textbf{k}_{\text{basal}}$ and $\textbf{k}_{\text{bolus}}$ values. For a single dose of basal insulin, $\textbf{x}_{\text{basal}}$, or bolus insulin, $\textbf{x}_{\text{bolus}}$, administered at time, $t$, we obtain the plasma insulin concentration at that time for patient, $i$:
@@
Treating each dose category as a distinct entity is an approximation and may not capture all nuances of dose-dependent kinetics. However, all physiological models approximate complex biological processes. Our goal is to develop models validated under established linear PK assumptions from the pharmacology literature while still accounting for dose-dependent effects, balancing simplicity and precision. This framework also ensures broader applicability across various medications, as most drugs are assumed to follow linear PK (Rosenbaum, 2017).
The equations presented are shown in a patient-specific manner to highlight that the PK encoder generates individualized information. However, these equations can be applied to produce patient-specific outputs for multiple patients simultaneously using vectorized operations.